OpenStudy (adi3):

f(x) = x^3 g(x) = x-4 Find the composite function: f(g(x))

OpenStudy (adi3):

@imqwerty

OpenStudy (shadowlegendx):

ԅ(≖‿≖ԅ) Learn math, you will ԅ(≖‿≖ԅ)

OpenStudy (adi3):

OMG

OpenStudy (anonymous):

i believe its function of function

OpenStudy (adi3):

??

OpenStudy (anonymous):

substitute in f(x) the x as x-4

OpenStudy (anonymous):

that's all :)

OpenStudy (adi3):



imqwerty (imqwerty):

f(x) = x^3 g(x) = x-4 f(g(x)) in f(x) just put x= g(x) that is equal to x-4

OpenStudy (anonymous):

yep

OpenStudy (adi3):

there is more

OpenStudy (anonymous):

to find f'(x)=?

OpenStudy (adi3):

the answer is x^3 + 4x^2 + 40x - 64

OpenStudy (anonymous):

first of all expand the function

OpenStudy (adi3):

thats the answer

OpenStudy (anonymous):

(a+b)^3=a^3+b^3+3a^2b+3ab^2

imqwerty (imqwerty):

so what

OpenStudy (adi3):

?? @secret88

OpenStudy (adi3):

@IrishBoy123 need help

OpenStudy (anonymous):

expand the function please..

OpenStudy (adi3):

x^3 - 64

OpenStudy (irishboy123):

this might help  you need to expand, but you wont get that answer.....

OpenStudy (anonymous):

hmm.. idk then that ans might be wrong

OpenStudy (adi3):

@secret88 could you help me expand

OpenStudy (anonymous):

( a – b )³ = a ³ – 3a² b + 3ab² – b³

OpenStudy (anonymous):

@Adi3 ?